OpenStudy (anonymous):

lim x --> 0 , ( 1 + (cotx)^2) /( cscx )^3

ganeshie8 (ganeshie8):

use your trig

OpenStudy (anonymous):

could i use l'hopital rule

ganeshie8 (ganeshie8):

i would try simplifying the expression first

OpenStudy (anonymous):

(2cotx (-(cscx)^2) )/ 3 cscx (-(cscxcotx) = 2/3, but the simplified one is equal to something else?

ganeshie8 (ganeshie8):

( 1 + (cotx)^2) /( cscx )^3 = (cscx)^2 / (cscx)^3 = 1/cscx = sinx

OpenStudy (anonymous):

so therefore i cannot use l'hopitals rule and must simplify instead?

ganeshie8 (ganeshie8):

you need to check if the expression is in indeterminate form before using l'hopitals

OpenStudy (anonymous):

i got 1 +infinity over +/-infinity is that allowed, so infinity/infinity form

ganeshie8 (ganeshie8):

trying lhopital here is just ridiculous here as it simplifies nicely without lhopital

OpenStudy (anonymous):

why is the answers different?

ganeshie8 (ganeshie8):

yeah im thinking the same, it should not fail if it is indeterminate

OpenStudy (anonymous):

you only can use lim sinx/x =1 or anything derived from it

OpenStudy (anonymous):

so this equation isn't derived from sinx/x? which is the reason it doesn't work

OpenStudy (anonymous):

well tricky a bit it would be something solvable lets say m lim m . x =0

OpenStudy (anonymous):

limit of a equation "m" as x approaches 0?

OpenStudy (anonymous):

i just dont wanna complicated u lol why dont u work on it simply and expand all terms ?

OpenStudy (anonymous):

limit 1/csc x is not infinity :O

ganeshie8 (ganeshie8):

limit of (cscx)^3 is not infinity as x->0

OpenStudy (anonymous):

 ? plus or minus infinity

OpenStudy (anonymous):

just work simply